For two given sets of data it is known that x bar is 10 and y bar is 4,the gradient of the regression line y on x is 0.6 .Find the equation of the regression line and estimate y when x = 2
1
Expert's answer
2020-02-26T11:04:24-0500
"\bar{y}=a+b\bar{x}"

"b=0.6"

Given that "\bar{y}=4, \bar{x}=10". Then


"4=a+0.6(10)"

Solve for "a"


"a=4-6""a=-2"

The equation of the regression line 


"\bar{y}=-2+0.6\cdot\bar{x}"

Estimate "y" when "x=2"


"y=-2+0.6\cdot2""y=-0.8"